OpenStudy (anonymous):

I need the steps to this answer: x-1/4 - x-1/6=1

OpenStudy (anonymous):

(x-1)/4 - (x-1)/6=1 3(x-1) - 2(x-1)=12 3x-3-2x+2=12 x-1=12 x=13 (13-1)/4 - (13-1)/6 = 1 12/4 - 12/6 = 1 3 - 2 = 1